# Detecting LLM-Written Peer Reviews

### v1

@@ 3.2.1 Defense strategy 1: paraphrasing
Table 4(a) provides the paraphrased accuracies. As expected, the random start watermarks do not remain in the paraphrased reviews, since this watermark describes a very specific sequence of words. Adding random citations to the review remain a higher percentage of the time when the review is paraphrased.
@@ 3.2.3 Defense strategy 3: crop out end of paper
Gemini 1.5 Pro 0.94 0.0 0.81
ChatGPT 4o 1.0 0.0 0.87
Claude 3.5 Sonnet 0.98 0.0 0.95
(a) Paraphrased accuracies for various LLMs. The Random Citation watermark refers to cases where reviews start with the random citation. Success is determined if the citation remains in any location.
@@
0.19 0.03 0.09
(b) Accuracies in identifying watermarks present in the paper using ChatGPT 4o with White Text Injection.
@@
0.46 0.21
(c) Accuracies when injecting the watermark into a page other than the last.
Table 4: Accuracies across various reviewer defenses
@@ 3.3 Cryptic prompt injection
In our experimental evaluation of the cryptic prompt injection, we use two white-box LLMs, Llama 2 [TMS+23] and Vicuna 1.5 [CLL+23], to generate reviews for 20 examples each from two datasets, International Congress on Peer Review and Scientific Publication (PRC) 2022 abstracts and PeerRead papers [KAD+18]. We execute the Greedy Coordinate Gradien (GCG) algorithm [ZWC+23] for 6000 iterations on an NVIDIA A100 or H100 GPU with 80 GB of memory. Due to computational limitations, we truncate the input prompt to 2000 tokens to avoid out-of-memory errors. For each LLM, we optimize the tokens of the cryptic prompt with respect to the specific model, using its gradients to update the tokens in each iteration. We set the number of optimizable tokens in the cryptic prompt to 30. See Appendix B for the hyperparameter settings of the GCG algorithm used in our experiments.
@@
Table 5: Performance of cryptic prompt injection for two LLMs, Llama 2 and Vicuna 1.5, and two datasets, PRC Abstracts and PeerRead Papers, at different iterations of the optimization algorithm.
@@ 3.4 Grant proposals
We also run our watermarking strategies on NSF Grant Proposals obtained from the Open Grants website [ope24]. There are 52 proposals in this dataset and we run them all on the three watermarking strategies described in Sec. 2.1 using white text injection. We use ChatGPT 4o API to generate the reviews and report the accuracies in Table 6. We observe the accuracies to be reasonably high.
@@
Table 8: Empirical evaluation of the False Positive Rate (FPR) and True Positive Rate (TPR) for individual reviews using Algorithm 1 and comparing it with the theoretical bound on the FPR obtained in Proposition 1. The value of $k$ is chosen here to upper bound the FPR at specified values. Note that the ICLR 2021 review set and the ICLR 2024 review set are augmented with 100 LLM-generated reviews with an embedded watermark.
@@ 3.5 Control condition
Table 9: Observing the False Positive Rate (FPR), True Positive Rate (TPR), the number of reviews discarded ($|\mathcal{I}|$), and the number of watermarks discarded ($|\mathcal{J}|$) to control the Family-Wise Error Rate $\alpha$ at various levels, using Algorithm 2 with Algorithm 3 as subroutine. Note that we report the results by using the default values of $\rho$ and $\Omega$ in Algorithm 3. The ICLR 2021 review set and the ICLR 2024 review set are augmented with 100 LLM-generated reviews with an embedded watermark. Observe that under FWER control, no reviews are falsely flagged. The random citation watermark has a high detection rate due to the large enough size of $\mathcal{W}$.
@@ 4 Discussion
We propose a method to identify reviews generated by large language models. [^6] On a somewhat related note, with the growing discussion around using LLMs as reviewers [LS23, LZC+23, DHBD24], it is important to recognize that authors can exploit indirect prompt injections in their PDFs to significantly bias LLM reviewers in a favorable manner. See Appendix C for an evaluation. We repurpose what is typically seen as a vulnerability—indirect prompt injection—for a beneficial application. We provide statistical detection approaches that can control the family-wise error rate, independent of potentially evolving human-writing styles, with a higher power than standard approaches. Our evaluations show that this approach maintains a high success rate, even when facing multiple reviewer countermeasures. For broader applications of watermarking in LLMs for general text [KGW+23, FCT+23], future research could explore adapting the ideas presented here — preprocessing of the text independently of random choices — to effectively control the family-wise error rate when evaluating multiple texts. As in any computer security application, detecting such illicit activities, and the efforts by bad actors to circumvent detection, remains an ever-evolving cat-and-mouse game. We hope that our work, along with contributions from the community, will continue to enhance the robustness of these detection techniques. Our watermarking and obfuscation schemes can be adapted into new variants; given the public disclosure of our current methods, future implementations may employ alternative variants while adhering to the core design principles we have outlined.
@@ Appendix A Examples of outputs of LLMs in various experimental settings
Table 10: Examples of Reviews from Random Start Watermark
@@
Table 11: Examples of Reviews from Random Technical Term Watermark
@@
Table 12: Here, we provide examples of reviews from the Random Citation watermark. While we primarily experimented on starting the review with a random fake reference, we also include some examples where we asked the LLM to include a random fake reference somewhere in the middle of the review.
@@
Table 13: Examples of Paraphrased Reviews Alongside Originals Generated by LLMs
@@ Appendix C Inducing LLMs to provide higher ratings in generated reviews
Without instruction 7.0±$\pm{0.26}$ 6.96±$\pm{0.27}$
Table 14: Mean ($\pm$ standard deviation) ratings with high-rating-inducing obfuscated instruction

### v2

@@ 3.2.1 Defense strategy 1: paraphrasing
Table 4(a) provides the paraphrased accuracies with 95% confidence intervals. As expected, the random start watermarks do not remain in the paraphrased reviews, since this watermark describes a very specific sequence of words. Adding random citations to the review remain a higher percentage of the time when the review is paraphrased.
@@ 3.2.3 Defense strategy 3: crop out end of paper
Table 4: Accuracies with 95% confidence intervals across various reviewer defenses using bootstrap resampling with 10,000 replicates.
(a) Accuracies with 95% confidence intervals in paraphrasing task for various LLMs. The random citation watermark refers to cases where reviews start with the random citation. For both the random citation and technical term watermark, the criterion for success is if the watermark remains in any location.
@@
Gemini 2.0 Flash 0.94 ± 0.05 0.0 ± 0.0 0.81 ± 0.08
ChatGPT 4o 1.0 ± 0.0 0.0 ± 0.0 0.87 ± 0.07
Claude 3.5 Sonnet 0.98 ± 0.02 0.0 ± 0.0 0.95 ± 0.05
(b) Accuracies with 95% confidence intervals in identifying watermarks present in the paper using ChatGPT 4o with White Text Injection.
@@
0.19 ± 0.09 0.03 ± 0.03 0.09 ± 0.06
(c) Accuracies with 95% confidence intervals when injecting the watermark into a page other than the last.
@@
0.46 ± 0.14 0.21 ± 0.11
@@ 3.3 Cryptic prompt injection
In our experimental evaluation of the cryptic prompt injection, we use two white-box LLMs, Llama 2 [TMS+23] and Vicuna 1.5 [CLL+23], to generate reviews for 20 examples each from two datasets, International Congress on Peer Review and Scientific Publication (PRC) 2022 abstracts and PeerRead papers [KAD+18]. We execute the Greedy Coordinate Gradien (GCG) algorithm [ZWC+23] for 6000 iterations on an NVIDIA A100 or H100 GPU with 80 GB of memory. Due to computational limitations, we truncate the input prompt to 2000 tokens to avoid out-of-memory errors. For each LLM, we optimize the tokens of the cryptic prompt with respect to the specific model, using its gradients to update the tokens in each iteration. We set the number of optimizable tokens in the cryptic prompt to 30. See Greedy coordinate gradient algorithm for the hyperparameter settings of the GCG algorithm used in our experiments.
@@
Table 5: Performance of cryptic prompt injection for two LLMs, Llama 2 and Vicuna 1.5, and two datasets, Peer Review Congress (PRC) Abstracts and PeerRead Papers, at different iterations of the optimization algorithm.
@@ 3.4 Grant proposals
We also run our watermarking strategies on NSF Grant Proposals obtained from the Open Grants website [ope24]. There are 52 proposals in this dataset and we run them all on the three watermarking strategies described in Section 2.1.1 using white text injection. We use ChatGPT 4o API to generate the reviews and report the accuracies with 95% confidence intervals in Table 6. We observe the accuracies to be reasonably high.
Table 6: Accuracies with 95% confidence intervals on grant proposals with GPT 4o using bootstrap resampling with 10,000 replicates.
@@
Table 8: Empirical evaluation of the False Positive Rate (FPR) and True Positive Rate (TPR) for individual reviews. Results obtained using Algorithm 1 and comparing it with the theoretical bound on the FPR obtained in Proposition 1. The value of $k$ is chosen here to upper bound the FPR at specified values. Note that the ICLR 2021 review set and the ICLR 2024 review set are augmented with 100 LLM-generated reviews with an embedded watermark.
@@ 3.5 Control condition
Table 9: False Positive Rate (FPR), True Positive Rate (TPR), the number of reviews discarded ($|\mathcal{I}|$), and the number of watermarks discarded ($|\mathcal{J}|$) to control the Family-Wise Error Rate $\alpha$ at various levels. Results obtained using Algorithm 2 with Algorithm 3 as subroutine. Note that we report the results by using the default values of $\rho$ and $\Omega$ in Algorithm 3. The ICLR 2021 review set and the ICLR 2024 review set are augmented with 100 LLM-generated reviews with an embedded watermark. Observe that under FWER control, no reviews are falsely flagged. The random citation watermark has a high detection rate due to the large enough size of $\mathcal{W}$.
@@ 4 Discussion
We propose a method for identifying peer reviews generated by large language models (LLMs), repurposing what is typically viewed as a vulnerability—indirect prompt injection—for a beneficial application. Our framework includes statistical detection techniques that offer formal control over the family-wise error rate (FWER), independent of the variability in human writing styles, and with higher power than standard correction methods. Our empirical evaluations show that the approach maintains a high watermarking success rate and strong detection accuracy, even when facing common reviewer defenses such as paraphrasing.
While our focus has been on detecting LLM-generated peer reviews, the underlying ideas can extend more broadly. Future research may explore adapting our watermarking and detection techniques to general text generation settings, building on prior work in LLM watermarking [KGW+23, FCT+23]. For example, preprocessing text independently of watermark selection may offer scalable and statistically sound ways to control FWER when evaluating multiple documents. As in other domains of computer security, detecting misuse—and the inevitable attempts to evade detection—remains an ongoing challenge. We hope our contributions, along with future work from the research community, will continue to enhance the robustness and reliability of detection strategies. Finally, with the growing interest in using LLMs as reviewers [LS23, LZC+23, DHBD24], it is important to acknowledge the risk that authors themselves could exploit indirect prompt injection to bias LLM-generated reviews in their favor. We present a preliminary evaluation of this threat in Inducing LLMs to provide higher ratings in generated reviews, highlighting the need for safeguards on both sides of the review process.
@@ Appendix A Examples of outputs of LLMs in various experimental settings
Table 10: Examples of reviews from random start watermark.
@@
Table 11: Examples of reviews from random technical term watermark.
@@
Table 12: Examples of reviews from the random citation watermark.
@@
While we primarily experimented on starting the review with a random fake reference, we also include some examples where we asked the LLM to include a random fake reference somewhere in the middle of the review.
Table 13: Examples of paraphrased reviews alongside originals generated by LLMs.
@@ Appendix C Inducing LLMs to provide higher ratings in generated reviews
Table 14: Mean ($\pm$ standard deviation) ratings with high-rating-inducing obfuscated instruction.